OpenStudy (anonymous):

what is 32^-3/5? need help understanding radicals

OpenStudy (mertsj):

\[a ^{-n}=\frac{1}{a^n}\]

OpenStudy (anonymous):

ok what now?

OpenStudy (mertsj):

\[32^{\frac{-3}{5}}=\frac{1}{32^{\frac{3}{5}}}=\frac{1}{2^3}=\frac{1}{8}\]

OpenStudy (anonymous):

ok